# Reblurring-Guided Single Image Defocus Deblurring: A Learning Framework with Misaligned Training Pairs

### v1

@@ 3.3.2 Reblurring Loss $\mathcal{L}_{R}$
${\mathcal{L}_{R}}=\sqrt{\parallel\bm{I}_{B}-\hat{\bm{I}}_{B})\parallel^{2}+\varepsilon^{2}},$ (12)
@@
3.4 Defocus Map Loss $\mathcal{L}_{P}$ with Pseudo Supervision
@@
where $\textit{Repmat}_{i}$ duplicates matrix $\bm{W}_{i}$ for $i$ times to ensure that kernel seeds $\bm{s}_{i}$ and their corresponding weighting matrices ${\bm{{W}}}_{i}$ have matching dimensions. Finally, during training of the prediction network, pseudo defocus map ${\bm{{B}}_{P}}$ serves as supervision
@@
3.5 A New Dataset for image defocus deblurring
To validate the effectiveness of our approach on a specific device, we employed a HUAWEI X2381-VG camera to establish a new dataset for image defocus deblurring, referred to SDD dataset. This camera has adjustable DOF through vertical lens movement and aperture modulation, allowing us to collect pairs of blurry images and corresponding ground-truth sharp images from the same scene. Despite our best efforts to ensure proper alignment between the blurry images and ground-truth sharp images, the SDD dataset exhibits noticeable misalignment in training pairs compared to DPDD, as depicted in Fig. 1. This discrepancy can be attributed to adjustments made by using an electronic motor to manipulate the camera lens.
@@ 4 Experiments
  • DPDD dataset was built by using a dual-pixel camera, capturing the defocus and all-in-focus pairs in two successive shots. It contains 350/76/74 image triplets for training/testing/validation, respectively. Each blurry image is paired with a corresponding sharp image. We use 350 blurry-sharp image pairs for training and 76 blurry images for testing.
@@
Evaluation Metrics: We report the three most widely used metrics to quantitatively assess the compared methods: Peak Signal-to-Noise Ratio (PSNR), Structural Similarity Index (SSIM)[44], and Learned Perceptual Image Patch Similarity (LPIPS)[45].
@@ 4.1 Implementation Details
At the beginning of training, the predicted optical flow $\bm{\Phi}_{\bm{I}_{S}\rightarrow\hat{\bm{I}}}$ and $\bm{\Phi}_{\hat{\bm{I}}\rightarrow\bm{I}_{S}}$ are hardly informative due to the low quality of $\hat{\bm{I}}$. Therefore, during the initial training stage, we calculate $\bm{\Phi}_{\bm{I}_{S}\rightarrow\hat{\bm{I}}}$ using $\mathcal{F}_{flow}(\bm{I}_{S},\bm{I}_{B})$ over $T$ training epochs, and subsequently by Eq. (4). To ensure the quality of $\hat{\bm{I}}$, we empirically set $T$ as 15 in our experiments. During training, we set $\lambda$ as 0.35 to generate the calibration masks and the maximal radius of blur kernels $m$ is set as 8. The trade-off parameters $\alpha$ and $\beta$ in Eq. (2) are both set as 0.5.
@@ 4.2 Evaluation on SDD Dataset
In this section, we compare our proposed defocus deblurring method against state-of-the-art single-image defocus deblurring approaches, including JNB [9], EBDB [48], DMENet [49] and DPDNetS [12]. Notably, the first three methods are primarily developed for defocus map estimation, we follow [10, 11] to get the final estimated sharp images using non-blind deconvolution. Additionally, we retrain some state-of-the-art deblurring methods on the SDD dataset for further comparison, including DMPHN  [50], MPRNet [51], Restormer[15], Loformer [22] and GRLS-B [18]. It is noteworthy that this study focuses on single-image defocus deblurring using misaligned data pairs. GRL [18], for instance, has a dual-pixel version, but only its single image version (GRLS-B [18]) is employed in our experiments. Since our dataset is designed for tackling misalignment tasks, directly calculating pixel-level metrics may not be accurate. Therefore, when testing all methods, we used an existing pretrained optical flow network to warp the sharp images $\bm{I}_{S}$, aligning them with the deblurred images $\hat{\bm{I}}$, and then calculated the evaluation metrics between $\hat{\bm{I}}$ and $\bm{I}_{S}^{w}$.
Table 1 summarizes the results of all competing methods on SDD dataset. In our conference paper [21], our method directly adopt MPRNet as the deblurring model, and achieves notable performance gains over MPRNet. In this work, our reblurring-guided learning frameworks Eq. (1) and Eq. (2) both obtain better deblurring performance, while Ours (Eq. (2)) further improves the quantitative metrics benefiting from the defocus map as degradation prior in the baseline deblurring network. Nevertheless, all the models trained by reblurring-guided framework can well handle the misalignment issues in training data. In the following, the term “Ours” refers to our proposed method described by Eq. (2), unless otherwise specified for clarity. The visualizations shown in Fig. 8 prove that our method overcomes the limitations of misaligned training pairs, achieving superior restoration of fine details in the presence of strong defocus blur. For example, in the second row of images, our method produces the sharpest lines while avoiding blur artifacts and distortions.
Table 1: Quantitative comparison of competing methods on SDD dataset. Ours [21] is the method proposed by us in the conference paper, while Ours (Eq. (1)) and Ours (Eq. (2)) are the methods proposed in this work.
@@
JNB [9] 24.06 0.720 0.443
EBDB [48] 24.06 0.720 0.442
DMENet [49] 24.02 0.732 0.426
    DMPHN [50]  25.00  0.769   0.326
DPDNetS [12] 24.81 0.760 0.343
MPRNet [51] 26.28 0.796 0.302
Restormer[15] 26.39 0.806 0.301
GRLS-B[18] 26.43 0.809 0.298
Loformer[22] 26.51 0.808 0.296
Ours [21] 26.88 0.810 0.265
Ours (Eq. (1)) 26.85 0.809 0.269
 Ours (Eq. (2)) 26.92 0.813 0.251
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Input DPDNetS [12] MPRNet [51] Restormer[15] Loformer[22] Ours GT
Figure 8: Visual comparison of competing methods on SDD dataset.
@@ 4.3 Evaluation on DPDD Dataset
For the DPDD dataset, besides the methods mentioned above, two state-of-the-art single image defocus deblurring methods Son et al. [24] and IFAN [14] are also compared. It is worth noting that although the DPDD dataset is meant to be aligned, there actually exists slight misalignment as shown in Fig. 1. Therefore, we calculate the evaluation metrics in view of $\hat{\bm{I}}$ v.s. $\bm{I}_{S}^{w}$. The experimental results are reported in Table 2.
Table 2: Quantitative comparison on DPDD dataset [12].
@@
JNB [9] 21.30 0.692 0.458
EBDB [48] 22.66 0.737 0.397
DMENet [49] 22.89 0.740 0.388
DPDNetS [12] 24.48 0.778 0.262
Son et al.[24] 25.49 0.807 0.212
IFAN [14] 25.86 0.825 0.192
MPRNet [51] 26.03 0.820 0.214
Restormer[15] 26.65 0.850 0.158
Loformer[22] 26.10 0.840 0.197
GRLS-B[18] 26.67 0.853 0.159
 Ours 26.75 0.852 0.155
From Table 2, we have a similar observation as in Table 1, i.e., the learning-based methods outperform the traditional ones, and the deblurring performance can be further boosted by adopting our reblurring-guided learning framework. The best performed metrics are highlighted in bold, while the second-best metrics are underlined. By tolerating the slight misalignment existing in DPDD dataset, JDRL contributes to better performance on testing set of DPDD. Notably, the DPDD dataset was collected under strictly controlled conditions using remote control and strictly controlled capture conditions to ensure precise alignment. However, in real-world scenarios, it is impossible to achieve perfectly aligned blurry-sharp pairs through manual control. Our misalignment training strategy can, to some extent, compensate for the discrepancies introduced during the data collection process.
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Input Son et al. [24] IFAN [14] Restormer[15] GRLS-B[18] Ours GT
Figure 9: Visual comparison of competing methods on RealDOF dataset. Ours performs visually better than the other methods in terms of textures and structures.

### v2

@@ 3.2.2 Reblurring Loss $\mathcal{L}_{R}$
${\mathcal{L}_{R}}=\sqrt{\parallel\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{\bm{I}_{B}-\hat{\bm{I}}_{B}}\parallel^{2}+\varepsilon^{2}},$ (7)
@@
The final issue lies in the acquisition of pseudo defocus map supervision $\bm{B}_{P}$. As shown in Eq. (6), a reblurred image can be considered as a weighted summation of blurred images generated by isotropic blur kernels. Consequently, the blur amount measurement at a pixel coordinate in the blurred image can be obtained through the utilization of weighted kernel seeds. Therefore, obtaining $\bm{B}_{P}$ with size $H\times W\times M$ involves assigning weights $\bm{W}$ with size $H\times W\times m$ to the kernel seeds $\bm{S}$ with size $H\times W\times M$.
3.2.3 Derivation of Pseudo Defocus Map
@@
where $\odot$ is element-wise multiplication, and Repmat duplicates matrix $\bm{W}_{i}$ for $i$ times to ensure that kernel seeds $\bm{s}_{i}$ and their corresponding weighting matrices ${\bm{{W}}}_{i}$ have matching dimensions. This process is shown in Fig. 6. Subsequently, pseudo defocus map $\bm{B}_{P}$ will be utilized to train the deblurring module as presented in Sec. 3.3.2, supervising the estimation of defocus blur map.
Figure 6: Visualization of Repmat: The matrix $\bm{W}_{i}$ is replicated $i$ times, enabling feasible element-wise multiplication with $\bm{S}\in\mathbb{R}^{H\times W\times M}$, where $M=\sum_{i=2}^{m}i$. Note that $\bm{W}_{1}$, which corresponds to the zero-blur-level image $\hat{\bm{I}}_{B}^{1}$, does not contribute to the blur amount, and thus is excluded before Repmat operation.
3.3 Deblurring Module
We present the deblurring network, and the training loss functions $\mathcal{L}_{D}$ and $\mathcal{L}_{P}$.
3.3.1 Deblurring Model
Upon existing deblurring network $\mathcal{D}$, e.g., the CNN-based or Transformer-based deblurring networks, we introduce a blur map estimator $\mathcal{E}$ and a fusion block $\mathcal{F}$, by which the estimated blur map $\hat{\bm{B}}=\mathcal{E}(\bm{I}_{B})$ and fused image $\bm{I}_{f}=\mathcal{F}(\bm{I}_{B},\hat{\bm{B}})$. Taking the blurry image $\bm{I}_{B}$ as input, the deblurring module can be formulated as
$\hat{\bm{I}}=\mathcal{D}(\bm{I}_{f})=\mathcal{D}\left(\mathcal{F}\left(\bm{I}_{B},\mathcal{E}({\bm{I}_{B}})\right)\right).$ (9)
Since we do not modify the architecture of $\mathcal{D}$, we in the following only focus on blur map estimator $\mathcal{E}$ and fussion block $\mathcal{F}$.
Defocus Blur Map Estimator $\mathcal{E}$: The defocus blur map estimator is designed to predict the defocus blur map, denoted as $\hat{\bm{B}}=\mathcal{E}(\bm{I}_{B})$. This lightweight network consists of 2 convolutional layers and 3 residual layers. However, since no direct supervision is available for the estimated defocus blur map $\hat{\bm{B}}$, it may result in significant deviation from the true defocus blur map. To address this issue, we utilize the pseudo ground-truth defocus maps generated by our reblurring module, as described in Sec. 3.2.3, and introduce a dedicated loss term $\mathcal{L}_{P}$ to optimize the estimator.
Fusion Block $\mathcal{F}$: Empirically, we observe that the defocus blur map and blurry image also have misalignment issues, especially for early training epochs. Therefore, we explore the fusion strategy by employing deformable attention mechanism as illustrated in Fig. 7 to effectively incorporate the degradation-related prior for image deblurring. The deformable attention mechanism allows the model to focus on relevant spatial regions of the pseudo defocus blur maps while dynamically adjusting the attention weights based on the local features of the blurred image.
Given the estimated defocus blur map $\hat{\bm{B}}\in\mathbb{R}^{H\times W\times M}$ and the blurry input image $\bm{I}_{B}\in\mathbb{R}^{H\times W\times 3}$, we introduce a deformable cross-attention mechanism to adaptively fuse information across different feature representations. For each pixel coordinate $(x,y)$ in the spatial domain, the fusion process of the deformable cross-attention is formulated as follows:
$\begin{split}\bm{I}_{f}(x,y)=&\bm{I}_{B}(x,y)\!+\!\bm{W}_{o}\sum_{n=1}^{N_{p}}\left(\bm{W}_{q(n)}\bm{I}_{B}(x,y)\right)\odot\\ &\left(\bm{W}_{v(n)}\hat{\bm{B}}{\left(x\!+\!\Delta x_{n},\ y\!+\!\Delta y_{n}\right)}\right),\end{split}$ (10)
where $\odot$ is element-wise multiplication, and $N_{p}$ is the number of deformable sampling points, which controls the spatial scope of attention. Specifically, $\bm{I}_{B}(x,y)\in\mathbb{R}^{3\times 1}$ denotes the feature vector of the input image $\bm{I}_{B}$ at position $(x,y)$, while $\hat{\bm{B}}{(x,y)}\in\mathbb{R}^{M\times 1}$ represents the corresponding feature from the blur map $\hat{\bm{B}}$ sampled at position $(x,y)$. The learnable weights $\bm{W}_{q(n)}\in\mathbb{R}^{M\times 3}$ and $\bm{W}_{v(n)}\in\mathbb{R}^{M\times M}$ are used to project features in blurry image and defocus blur map, respectively. $\Delta x_{n}$ and $\Delta y_{n}$ are the spatial offsets. $\bm{W}_{o}\in\mathbb{R}^{3\times M}$ is the output projection matrix that maps the aggregated features back to the original image dimension. The fused image $\bm{I}_{f}\in\mathbb{R}^{H\times W\times 3}$ can then be fed to existing deblurring network without modifying on architecture. This strategy overcomes the limitations of traditional fixed attention mechanisms by enabling more flexible integration of information from defocus blur maps, leading to better alignment of the blur prior with the image content.
Figure 7: Overview of our fusion block for integrating defocus map and input blurry image, which incorporates a deformable attention mechanism that better aligns the blur prior with the image content. The fused image $\bm{I}_{f}$ is of size $H\times W\times 3$ that can be fed to existing deblurring networks without architecture modification.
The architecture of $\mathcal{F}$ employs 5 attention heads, with each head using 4 sampling points. Each attention head independently computes its attention weights and feature fusion process, and the outputs from all heads are subsequently merged that enables the model to focus on relevant spatial regions of the pseudo defocus blur maps. The parameters of $\mathcal{D}$, $\mathcal{E}$ and $\mathcal{F}$ can be learned by optimizing a single deblurring loss $\mathcal{L}_{D}$. Moreover, to supervise the estimated defocus blur map $\hat{\bm{B}}$, we also introduce the dedicated loss $\mathcal{L}_{P}$.
[…]
Figure 8: Errors in optical flow estimation can incorrectly deform the sharp ground-truth image, e.g., in areas marked by green rectangles. Calibration masks can help by filtering out this adverse region.
3.3.2 Deblurring Losses $\mathcal{L}_{D}$ and $\mathcal{L}_{P}$
where $\mathcal{W}$ denotes linear warping operation [43].
To address the significant deviation of the estimated defocus blur map $\hat{\bm{B}}$ from the true defocus blur map, we introduce an additional loss term $\mathcal{L}_{P}$ that uses the pseudo supervision $\bm{B}_{P}$ derived in Section 3.2.3 as supervision
@@ 3.3.2 Deblurring Losses $\mathcal{L}_{D}$ and $\mathcal{L}_{P}$
3.4 A New Dataset for Image Defocus Deblurring
To validate the effectiveness of our approach on a specific device, we employed a HUAWEI X2381-VG camera to establish a new dataset for image defocus deblurring, referred to SDD dataset. This camera has adjustable DOF through vertical lens movement and aperture modulation, allowing us to collect pairs of blurry images and corresponding ground-truth sharp images from the same scene. Despite our best efforts to ensure proper alignment between the blurry images and ground-truth sharp images, the SDD dataset exhibits noticeable misalignment in training pairs compared to DPDD [1], as depicted in Fig. 1. This discrepancy can be attributed to adjustments made by using an electronic motor to manipulate the camera lens.
@@ 4 Experiments
  • DPDD dataset was built by using a dual-pixel camera, capturing the defocus and all-in-focus pairs in two successive shots. It contains 350/76/74 image triplets for training/testing/validation, respectively. Each blurry image is paired with a corresponding sharp image. We use 7,000 processed blurry-sharp pairs for training and 76 blurry images for testing.
@@
Evaluation Metrics: We report five metrics to quantitatively assess the compared methods: Peak Signal-to-Noise Ratio (PSNR), Structural Similarity Index (SSIM)[44], Learned Perceptual Image Patch Similarity (LPIPS)[45], Fréchet Inception Distance (FID)[46] and Deep Image Structure and Texture Similarity (DISTS)[47]. In defocusing deblurring scenarios with misaligned training data, existing methods may restore the image, but they often lead to severe distortions and truncations. Quantitative metrics, e.g., PSNR and SSIM, are sensitive to pixel-level differences and may not fully capture the local structural improvements, such as edge sharpness and texture clarity, which are more perceptually significant. This aligns with the findings in image restoration research [47, 46], where perceptual quality is increasingly valued over strict pixel-level fidelity. The study [47] also analyzes the unsatisfactory performance of existing metrics on certain image restoration techniques, partially due to their low tolerance to spatial misalignment. Thus, in addition to PSNR and SSIM [44], we also calculated LPIPS [45], FID [46], and DISTS [47].
@@ 4.1 Implementation Details
At the beginning of training, the predicted optical flow $\bm{\Phi}_{\bm{I}_{S}\rightarrow\hat{\bm{I}}}$ and $\bm{\Phi}_{\hat{\bm{I}}\rightarrow\bm{I}_{S}}$ are hardly informative due to the low quality of $\hat{\bm{I}}$. Therefore, during the initial training stage, we calculate $\bm{\Phi}_{\bm{I}_{S}\rightarrow\hat{\bm{I}}}$ using $\mathcal{F}_{flow}(\bm{I}_{S},\bm{I}_{B})$ over $T$ training epochs, and subsequently by Eq. (11). To ensure the quality of $\hat{\bm{I}}$, we empirically set $T$ as 15 in our experiments. During training, we set $\lambda$ as 0.35 to generate the calibration masks. The maximal radius of blur kernels $m$ is set as 8, and the sampling points number $N_{p}$ in deformable attention is set as 4. The trade-off parameters $\alpha$ and $\beta$ in Eq. (2) are both set as 0.5.
@@ 4.2 Evaluation on SDD Dataset
In this section, we compare our proposed defocus deblurring method against state-of-the-art single-image defocus deblurring approaches. we retrain state-of-the-art deblurring methods on the SDD dataset for further comparison, including DMPHN  [50], MPRNet [51], UformerT [16], Restormer [15] and Loformer [22]. To validate the effectiveness of our method, We have applied the proposed framework to both CNN-based and Restormer-based deblurring methods. As shown in Table 1, our proposed framework is evaluated from two perspectives: (i) reblurring-based learning framework, i.e., deblurring model+Ours(Eq. (1)), and (ii) reblurring-based blurring framework and deblurring blocks $\mathcal{E}$ and $\mathcal{F}$, i.e., deblurring model+Ours(Eq. (2)). One can see that for plain UNet, PSNR gains +1.2dB and +1.58dB are obtained by Ours(Eq. (1)) and Ours(Eq. (2)), respectively. For Loformer with top performance, PSNR gains +0.44dB and +0.68dB are obtained by Ours(Eq. (1)) and Ours(Eq. (2)), respectively. Considering that UNet, Restormer and Loformer are representative network architectures, covering from plain CNN to top Transformer, we believe that our proposed framework can be successfully applied to improve existing deblurring models.
Since our dataset is designed for tackling misalignment tasks, directly calculating pixel-level metrics may not be accurate. Therefore, when testing all methods, we used an existing pretrained optical flow network to warp the sharp images $\bm{I}_{S}$, aligning them with the deblurred images $\hat{\bm{I}}$, and then calculated the evaluation metrics between $\hat{\bm{I}}$ and $\bm{I}_{S}^{w}$.
Table 1 summarizes the results of all competing methods on SDD dataset. In our conference paper [21], our method directly adopt MPRNet as the deblurring model, and achieves notable performance gains over MPRNet. In this work, our reblurring-guided learning frameworks Eq. (1) and Eq. (2) both obtain better deblurring performance, while Ours (Eq. (2)) further improves the quantitative metrics benefiting from the defocus map as degradation prior in the baseline deblurring network. Nevertheless, all the models trained by reblurring-guided framework can well handle the misalignment issues in training data.
@@
The visualizations shown in Fig. 10 prove that our method overcomes the limitations of misaligned training pairs, achieving superior restoration of fine details in the presence of strong defocus blur. In the green box in the fifth row, pixel misalignment leads to optimization discrepancies during training, causing severe distortion in the image restoration process of basic models such as DPDNet[1], MPRNet[51], and Restormer[15]. The edges of the table and chair, which should be straight, are restored into pronounced curves, which contradicts our visual perception. In the first row, a similar phenomenon can also be observed. Although Loformer[22] performs well in image restoration, it experiences local “collapse” in training tasks with misaligned data. In the green box, Loformer[22] exhibits a “truncation” phenomenon during the restoration process, resulting in poor visual quality. It is clear that this issue can be effectively addressed by our training framework.
[…]
UNet 24.62 0.758 0.344 81.82 0.212
DMPHN [50] 25.00 0.769 0.326 71.47 0.208
DPDNetS [1] 24.81 0.760 0.343 75.66 0.210
MPRNet [51] 26.28 0.796 0.302 62.32 0.202
UFormerT [16] 25.68 0.774 0.321 70.24 0.208
Restormer[15] 26.39 0.806 0.301 58.90 0.189
Loformer[22] 26.51 0.808 0.296 54.80 0.179
UNet* 25.82 0.783 0.305 58.53 0.181
MPRNet* 26.88 0.810 0.265 55.30 0.177
UformerT* 26.53 0.808 0.297 56.18 0.180
Restormer* 26.89 0.817 0.257 51.22 0.168
Loformer* 26.95 0.820 0.255 48.77 0.165
UNet 26.20 0.796 0.298 55.49 0.180
Restormer 27.08 0.819 0.253 50.64 0.166
Loformer 27.20 0.826 0.251 47.91 0.164
[…]
Method FLOPs(G) Params(M)
MPRNet [51] 6830 20.1
UFormerT [16] 42.7 5.20
Restormer[15] 564 26.1
Loformer[22] 331 27.9
MPRNet* 6830 20.1
UformerT* 42.7 5.20
Restormer* 564 26.1
Loformer* 331 27.9
$\text{Restormer}^{\dagger}$ 582 26.2
$\text{Loformer}^{\dagger}$ 348 28.0
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Input DPDNetS [1] MPRNet [51] Restormer[15] Loformer[22] Loformer* $\text{Loformer}^{\dagger}$ GT
Figure 10: Visual comparison of competing methods on SDD dataset.Loformer is trained using Ours(Eq. (1)),while $\text{Loformer}^{\dagger}$ is trained using ours(Eq. (2)).
@@ 4.3 Evaluation on DPDD Dataset
For the DPDD [1] dataset, besides the methods mentioned above, two state-of-the-art single image defocus deblurring methods Son et al. [24] and IFAN [2] are also compared. It is worth noting that although the DPDD dataset is meant to be aligned, there actually exists slight misalignment as shown in Fig. 1. Therefore, we calculate the evaluation metrics in view of $\hat{\bm{I}}$ v.s. $\bm{I}_{S}^{w}$. The experimental results are reported in Table 3.
Table 3: Quantitative comparison on DPDD dataset [1]. We only apply our framework on Restormer that achieves top performance on this dataset, resulting in Restormer.
@@
DPDNetS [1] 24.48 0.778 0.262 76.46 0.158
Son et al.[24] 25.49 0.807 0.212 60.93 0.135
IFAN [2] 25.86 0.825 0.192 52.79 0.133
MPRNet [51] 26.03 0.820 0.214 60.68 0.140
Restormer[15] 26.65 0.850 0.158 45.57 0.104
Loformer[22] 26.10 0.840 0.197 53.24 0.126
$\text{Restormer}^{\dagger}$ 26.81 0.854 0.152 43.22 0.102
By tolerating the slight misalignment existing in DPDD dataset [1], JDRL contributes to better performance on testing set of DPDD. Notably, the DPDD dataset was collected under strictly controlled conditions using remote control and strictly controlled capture conditions to ensure precise alignment. However, in real-world scenarios, it is impossible to achieve perfectly aligned blurry-sharp pairs through manual control. Our misalignment training strategy can, to some extent, compensate for the discrepancies introduced during the data collection process.
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Input Son et al. [24] IFAN [2] Restormer[15] $\text{Restormer}^{\dagger}$ GT
Figure 11: Visual comparison of competing methods on the DPDD dataset and the RealDOF dataset. The first three rows are from the DPDD dataset, and the last three rows are from the RealDOF dataset. Our method demonstrates better visual results in terms of textures and structures.